OpenStudy (anonymous):

ok a nice one Find all Pairs \((m,n)\) of positive integers such that\[\frac{n^2+1}{mn-1}\]is an integer too.

OpenStudy (anonymous):

m=2, n=3 , ratio=2 n=2, m=3, ratio=1

OpenStudy (anonymous):

m=1,n=2 m=1,n=3

OpenStudy (anonymous):

still missing some of solutions

OpenStudy (asnaseer):

I got these: m=1, n=2, ratio=5 m=1, n=3, ratio=5 m=2, n=1, ratio=2 m=2, n=3, ratio=2 m=3, n=1, ratio=1 m=3, n=2, ratio=1

OpenStudy (anonymous):

quite right

OpenStudy (asnaseer):

I used a method where I assumed m = n + p and drew conclusions from the resulting equations

OpenStudy (asnaseer):

integer solutions could only exist if p=1, 2, -1 or -2

OpenStudy (asnaseer):

I'd be interested to know if there is another way of doing this

OpenStudy (anonymous):

i'd like also to see ur reasoning @asnaseer

OpenStudy (asnaseer):

sure - let me type it up for you...

OpenStudy (asnaseer):

let m = n + p where p is some other integer (negative, zero or positive), then we have:\[\frac{n^2+1}{mn-1}=\frac{n^2+1}{(n+p)n-1}=\frac{n^2+1}{n^2+np-1}\]for this expression to yield an integer, we must satisfy at least this:\[n^2+1\ge n^2+np-1\]\[\therefore1\ge np-1\]\[\therefore2\ge np\]\[\therefore n\le\frac{2}{p}\]

OpenStudy (asnaseer):

if p is negative, then this inequality becomes:\[n\gt\frac{2}{p}\]which then gives us:\[p=2\implies n\le1\implies n=1\]\[p=1\implies n\le2\implies n=1\text{ or }2\]\[p=-1\implies n\gt-2\implies n=1,2,3...,\infty\]\[p=-2\implies n\gt-1\implies n=1,2,3,...,\infty\]

OpenStudy (asnaseer):

I then took each case and created a table to see which combinations gave a valid integer ratio

OpenStudy (asnaseer):

e.g. p=2, n=1, m=n+p=1+2=3, ratio=1

OpenStudy (asnaseer):

hope there isn't a flaw in my reasoning?

OpenStudy (anonymous):

i enjoyed seeing ur solution...

OpenStudy (asnaseer):

thx :)

OpenStudy (asnaseer):

do you have an alternative method?

OpenStudy (anonymous):

yes its a little bit longer than this.

OpenStudy (asnaseer):

ok - I won't push you to post it, but if you could (even a scan of paper written solution) then I would really appreciate it.

OpenStudy (anonymous):

sure...i'll post it later. :) Nice to see this group alive again.

OpenStudy (asnaseer):

thx - and yes - finally it awakes! :)

OpenStudy (anonymous):

{m,n,ratio} {{1, 2, 5}, {1, 3, 5}, {2, 1, 2}, {2, 3, 2}, {3, 1, 1}, {3, 2, 1}}

OpenStudy (anonymous):

just a neat pint about this problem if \[mn-1|n^2+1\]so\[mn-1|m^2n^2-1+n^2+1=n^2(1+m^2)\]\[mn-1|1+m^2\]so if \((m,n)\) is answer \((n,m)\) will be answer

OpenStudy (anonymous):

*point

OpenStudy (anonymous):

n=1 gives m=2,3 so 4 answers from here : (1,2),(1,3),(2,1),(3,1) no answer for m=n suppose m>n\[kn-1=\frac{n^2+1}{mn-1}<\frac{n^2+1}{n^2-1}=1+\frac{2}{n^2-1}<2\]\[kn<3\]\[n=2 , k=1\]gives m=3 so 2 solution from here (2,3),(3,2)

OpenStudy (anonymous):

but why \[\frac{n^2+1}{mn-1}=kn-1\]\[\frac{n^2+1}{mn-1}=r\]its easy to show that\[r\equiv-1 \ \ \text{mod} \ n\]